OpenStudy (gabylovesyou):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1. −30 0 10 12

OpenStudy (gabylovesyou):

i got -12 ._.

OpenStudy (anonymous):

Nope I think..

OpenStudy (anonymous):

What is f(g(x)) tell me first..

OpenStudy (gabylovesyou):

plug in the g(x) equation to every x in the f(x) equation

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

\[f(g(x)) = 5(g(x)) -10\]

OpenStudy (gabylovesyou):

so f(-1 + 3) = 5 * -1 + 3 - 10

OpenStudy (anonymous):

\[= 5(x+3) - 10 = 5x + 5\]

OpenStudy (anonymous):

Keep left hand side as it is..

OpenStudy (gabylovesyou):

why 5x + 5 ..

OpenStudy (anonymous):

\[f(g(x)) =5x + 15 - 10 = 5x + 5\]

OpenStudy (gabylovesyou):

where did u get the 15 tho ?

OpenStudy (anonymous):